Suppose that the following equations characterize the economy;

C=$160+0.80Y_D I=$300 G=$200 T=$200

Please answer the following questions;

Write down the equilibrium condition (please use Z for the total demand for goods and Y for production).

Write down an equation for Z.

Calculate GDP (Y=?)

Calculate the multiplier

Calculate consumption?

What is saving?

Suppose that private firms are less confident about the future path of the economy. Firms cut investment by $100. Now the economy is experiencing an economic crisis. Calculate the new equilibrium GDP level.

Explain how the economy adjust to the new equilibrium. Please graph your answer.

Suppose that the government plans to guide the economy to recover from the crisis, however public are concerned about budget balance. Can the government help the economy without creating budget deficit? Support your answer by algebra.

Suppose that the following equations characterize the economy; C=c_0+c_1 Y_D=$160+0.80Y_D I=I ̅=$300 G=$200 T=t_0+t_1 Y=$50+0.10Y here t_0 is a lump-sum tax in which everybody pays the same amount of money. t_1 is the income tax rate. Please answer the following questions;

Write down an equation for Z Z=C+I+G Z=c_0+c_1 (Y-t_0-t_1 Y)+I+G 〖Z=c〗_0+c_1 Y-c_1 t_0-〖c_1 t〗_1 Y+I+G 〖Z=c〗_0+c_1 (1-t_1)Y-c_1 t_0+I+G Calculate GDP (Y=?) Y=Z Y〖=c〗_0+c_1 (1-t_1)Y-c_1 t_0+I+G Y=$160+0.80(1-0.10)Y-0.80*$50+$300+$200 Y=2,214.2

What is the multiplier in this economy? Y=Z 〖Y=c〗_0+c_1 (1-t_1)Y-c_1 t_0+I+G Y-c_1 (1-t_1)Y=c_0-c_1 t_0+I+G Y(1-c_1 (1-t_1))=c_0-c_1 t_0+I+G Y=1/(1-c_1 (1-t_1)) (c_0-c_1 t_0+I+G) or Y=1/((1-c_1+〖c_1 t〗_1)) (c_0-c_1 t_0+I+G) The multiplier=1/(1-c_1 (1-t_1))=1/((1-c_1+〖c_1 t〗_1))

What is the relationship between t_1 and the multiplier? What is the intuition for your answer? Calculate the slope of the total demand (ZZ) curve. Explain what automatic stabilizers mean?
Answer #1

(1)

In equilibrium, aggregate demand (Z)= Aggregate production (Y)

(2)

Z = C + I + G + NX

Y = 160 + 0.8(Y - T) + 300 + 200

Y = 660 + 0.8(Y - 200)

Y = 660 + 0.8Y - 160

0.2Y = 500

Y = 2,500

(3)

Multiplier = 1 / (1 - MPC) = 1 / (1 - 0.8) = 1 / 0.2 = 5

(4)

Consumption = 160 + 0.8 x (2,500 - 200) = 160 + 0.8 x 2,300 = 160 + 1,840 = 2,000

(5)

Saving = Y - C = 2,500 - 2,000 = 500

NOTE: As HOMEWORKLIB Answering Policy, 1st 5 parts have been answered.